OpenStudy (anonymous):

Calculate the discriminant. x2 – x + 2 = 0 A. 9 B. –7 C. –9 D. 7

OpenStudy (anonymous):

You have a quadratic equation. Use the quadratic formula: $$x=\frac{-b\pm\sqrt{ b^2 - 4ac}}{2a}$$ The discriminant is: $$\sqrt{ b^2 - 4ac}$$